# MICL: Improving In-Context Learning through Multiple-Label Words in Demonstration

### v1

@@ Abstract
In-context learning (ICL) enables large language models (LLMs) to perform new tasks by using sample-label pairs as demonstrations. However, variations in demonstrations can lead to significantly different performances. Current research mainly focuses on selecting demonstration samples, preassuming the class name to be the label word when creating sample-label pairs. However, the choice of label words is crucial for ICL performance. In addition, we observe that using a single class name in demonstration may not yield optimal results. In this paper, we propose to use multiple label words in one sample-label pair to enhance ICL performance. Further, we select and order sample-label pairs based on LLM’s output distribution, aiming to optimize the demonstration examples from both the samples’ and labels’ perspectives. Evaluation results on seven classification datasets show that the use of multiple label words, strategically organized by their selection, order and quantity, improves ICL performance through diverse label information.
MICL: Improving In-Context Learning through Multiple-Label Words in Demonstration
@@ 1 Introduction
In-context learning (ICL) could perform new tasks by using sample-label pairs from training data as demonstrations, without having to re-train or fine-tune large language models (LLMs) (Brown et al., 2020). The choice of demonstrations is crucial, as ICL performance can vary significantly with different organizations [^1] In this paper, we follow Wu et al. (2023) to denote the selection and ranking of sample-label pairs as organization. (Liu et al., 2022). To enhance ICL performance, most studies focus on the selection and ranking of samples in demonstrations (Zhang et al., 2022; Hongjin et al., 2022; Levy et al., 2023; Lu et al., 2022), preassuming the class name to be the label word, overlooking the importance of label word selection in demonstrations.
Label words in sample-label pairs may greatly impact ICL performance (Yoo et al., 2022). Fig. 1(a) shows 1-shot ICL performance under varied label words on five datasets, indicating that carefully selected label words in demonstrations could excel in both accuracy and robustness of ICL.
(a) Label effectiveness in ICL
(b) Label words output distribution in ZSL (more word evaluations are in Appx. A.2)
(c) Multiple label words effectiveness in ICL
@@
Figure 1: Exploration of label words in Llama2-7b. (a): We evaluate four sets of label words under the same samples across five datasets. Long bars indicate instability between seeds, while jumps between bars show accuracy differences among label sets. (b): We evaluate the logit values (y-axis) of various label words on SST-2 samples (x-axis) under zero-shot learning, showcasing 100 negative and positive samples per label word to demonstrate the logit separability of samples to label words. (c): In 1-shot ICL using class names, label-related words, and multiple label words (combining the two sets with spaces) as labels, performance with multiple label words surpassed the other two sets. (a), (b), (c) detailed analysis and experimental settings, including those on GPT2-xl, are in Appx. A.
LLM’s output distribution (logit) over labels is a key consideration in demonstration selection Rubin et al. (2022) and class prediction Wang et al. (2023). However, we find that certain class-related words, named as label words in this paper, may fit the LLM better than predefined class names. Fig. 1(b) shows the logit distribution of the label word "bad" and the class name "negative" under the zero-shot setting across negative and positive samples in a sentiment analysis dataset. Obviously, the label word "bad" exhibits a greater difference in logit for samples of the positive and negative classes than the class name "negative". This suggests that the commonly used class names may not be the best choice for label words in sample-label pair demonstrations for ICL.
The logits of label words vary significantly across samples. We found that only one label word may be insufficient to express the semantics of the class name. Expanding from a single class name to multiple label words, at the linguistic level, can reduce ambiguity and enrich the semantics of the label name, potentially leading to improved performance. As shown in Fig. 1(c), the use of multiple label words in sample-label pairs indeed improves ICL performance.
Verbalizer, which links words related to class names to label space, is beneficial for prompt-based learning Gao et al. (2021). However, directly employing all the label words in the verbalizer for ICL is infeasible. Firstly, verbalizer often contains hundreds of label words Hu et al. (2022), and inserting all these label words into ICL prompts can overwhelm the model or exceed token limits. Secondly, ICL relies on demonstration organization. Predefined label words from verbalizers may not fit well with the flexible nature of ICL. Without explicit instructions on sample pairing and order, improper sample-label pairs could mislead the model’s understanding. Besides, the impact of employing additional label words in prediction within ICL differs from that in prompt-based methods (Sec. 4.4).
In this work, we propose a new algorithm that selects and orders multiple label words for sample-label pairing based on LLM’s output distribution (logit) over training samples. We first filter the related label words collected from the large knowledge base to ensure they are tailored for LLM and dataset under study. We then apply zero-shot learning to training samples to obtain the LLM’s output distribution (logit) of the label words. We initialize the label in sample-label pairs using the word with the highest logit value, and then iteratively select and add extra label words to the sample-label pairs. The number of label words added to the sample-label pairs is determined in terms of ICL performance. To improve demonstration organization, we further select and rank samples based on the output distribution of their semantically-related label words and design their corresponding multiple label words.
@@
1. We propose MICL, a method that uses Multiple label words to enhance ICL. We develop an algorithm to filter related label words via samples’ output distribution, aiming to find label words that best suit the LLM and the data. By using multiple label words in sample-label pairing, more comprehensive label information is provided for ICL, which improves clarity and reduces ambiguity, and this in turn enhances ICL performance.
2. Based on the selected label words’ output distribution of training samples, we developed sample organization algorithm, involving selection and ordering of samples for demonstrations, which further improves ICL performance.
3. Extensive experiment results across various classification datasets prove MICL’s effectiveness.
@@ 2 Related Work
Figure 2: Overall architecture of MICL: Orange lines show Label Words Filtering workflow, blue lines represent Demonstration Sample Organization, and green lines depict Multiple-label Word Insertion.
@@
Label words matter in ICL The significance of label words in ICL has been debated. Yoo et al. (2022) demonstrates a positive relationship between ICL performance and accurate sample-label mapping. Li and Qiu (2023) reveals that using the same sample but different labels in the demonstrations can result in very different ICL performance. Wang et al. (2023) suggests that label words derive semantic representations from demonstrations for use in deep layers to make final predictions. Yu and Ananiadou (2024) further shows that these demonstration features are integrated into corresponding labels, with each in-context head extracting features specific to these labels. Milios et al. (2023) uses different label words in augmented samples to enhance ICL performance, indicating the potential for enhanced demonstration effectiveness when multiple label words are used. However, this approach significantly increases the required token length and running time in augmentation settings.
@@ 3 Method
In this section, we introduce our proposed method MICL, which comprises three parts: label words filtering, demonstration sample organization, and multiple-label word insertion, as shown in Fig. 2.
@@ 3.1 Problem Statement
Give a large language model $M$, class names (label space) $L$, a label words set $S$ with words related to class names, test sample $x_{test}$ and demonstrations $\left\{x_{i},y_{i}\right\}_{i=1}^{L}$ [^2] Our work focus on ICL with label-balanced demonstrations..
The zero-shot classification of $x_{test}$ can be based on the logits of class names only as
$\arg\max_{y\in\{L\}}p_{M}(y\mid x_{\text{test}})$ (1)
or based on the logits of class names and all label words:
$\arg\max_{y\in\{L,S\}}p_{M}(y\mid x_{\text{test}})$ (2)
Similarly, the 1-shot prediction of $x_{test}$ is
$\arg\max_{y\in\{L\}}p_{M}(y\mid x_{1}\oplus y_{1},\cdots,x_{L}\oplus y_{L}\oplus x_{\text{test}})$ (3)
or
$\arg\max_{y\in\{L,S\}}p_{M}(y\mid x_{1}\oplus y_{1},\cdots,x_{L}\oplus y_{L}\oplus x_{\text{test}})$ (4)
where $\oplus$ is the concatenation operation. The function $p_{M}(\cdot)$ returns the logits of words in $M$’s vocabulary. In the following sections, the prediction based on the logits of class names $L$ is referred to as class-name result, and the prediction based on the logits of all the label words $S$ and class names $L$ is referred to as label-words result.
As analyzed in Section 1, the use of class names only in the sample-label pair $x_{i}\oplus y_{i}$ may not be sufficient. In this study, we propose to use multiple label words $S_{i}$ to create sample-label pair $x_{i}\oplus S_{i}$, where $S_{i}$ is a sequence of multiple label words for class $i$. The selection and ordering of label words in $S_{i}$ will be introduced next.
3.2 Label Words Filtering
Given a set of label words relating to a class name, we developed a two-stage filtering algorithm to refine words, tailoring them for the LLM employed and the dataset under study.
Stage 1: Filtering Based on Separability of Logit Distribution Stage 1 filters out words based on the LLM’s logit distribution (Alg. 1, lines 3-14). We evaluate each label word’s separability based on its logit value across the training set under the zero-shot setting, where the input samples are formatted in a template without labels. The label word filtering is based on the following two principles: (1) label words whose logit values are not the highest for their own class’s training samples are discarded (Alg. 1, lines 7-9). These label words do not fit the samples of their own class, providing confusing information and negatively impacting the performance of ICL. (2) Retained words must have the maximum average non-negative logit values (Alg. 1, lines 10-12) to ensure semantic representativeness in LLM’s feedback. We observed that principle (1) is met across all datasets. Principle (2) is not satisfied in some datasets, so only principle (1) is applied to these datasets in the experiments.
Stage 2: Filtering Based on Point-Biserial Testing In Stage 2, the retained label words are further evaluated by Point-Biserial correlation testing Tate (1954). For a label word, if its mean logit value over training samples of its own class differs significantly from that of other classes, it will be retained, otherwise, it will be deleted from the label words set (Alg. 1, lines 15-22). The input in the testing is the logit vector $B\in\mathbb{R}^{1\times|D^{l}|}$ obtained from 0-shot learning in $D^{l}$ (Alg. 1, line 4). The label of $B$ is 0 if $l$ differs from the word’s label or 1 if they are the same. Significance testing is employed on the correlation results to ensure reliability.
Algorithm 1 Label Words Filtering
1:Input: Label set $S$, class names $L$, Train set $D$
2:Output: Refined label set $S_{r}$
3:Step 1: Distribution Separability Filtering
4:Perform zero-shot learning on $D$, split into sub-train sets $D^{l}$ for $l\in L$
5:for each word $w\in S$ with label $l\in L$ do
6:     Obtain average logit $b_{l}$ for $w$ in each $D^{l}$
7:     if label in $\max(\bigcup_{l\in L}b_{l})\neq l$ then
8:         Filter out $w$
9:     end if
10:     if $\max(\bigcup_{l\in L}b_{l})<0$ then
11:         Filter out $w$
12:     end if
13:end for
14:Return Refined label set $S_{1}$
15:Step 2: Point-Biserial Testing Filtering
16:for each word $w\in S_{1}$ do
17:     Compute Point-Biserial correlation $r$ for logit vector $B$ of $w$ in $D$
18:     if $r<0$ or ($r>0$ and $p$-value $>0.05$then
19:         Filter out $w$
20:     end if
21:end for
22:Return Refined label set $S_{r}$
3.3 Demonstration Sample Organization
After filtering, the remaining label words are semantically aligned with the dataset and LLM. Next, we solve the sample organization problem, including the selection of samples from each class and ordering of the samples in the demonstrations. Sample organization is based on the training samples’ LLM output distribution (logit) over the remaining label words $S_{r}$. For each sample (prompted in a template without label) in zero-shot learning, the LLM outputs the logit values for all words in $S_{r}$ obtained in Section 3.2. We rank the words based on their logits. We denote the number of words in $S_{r}$ belonging to class $l$ as $N_{l}$. With two requirements for the selected samples: (1) the words with correct sample-label semantic mappings have higher logits, and (2) to have these words ranked as high as possible within the top $N_{l}$, we employ two scoring methods, depending on how well the dataset meets the filtering principles.
Scoring method for datasets meeting both principles For sample $t_{j}$ in training set $D$ with label $l$, its score as a demonstration sample is the sum of the top-$N_{l}$ logit value in $S_{r}$ (Eq 5). If word $w_{i}$ does not belong to class $l$, we set logit $b_{w_{i}}=0$.
$score_{t_{j}}=\sum_{i=0}^{N_{l}}(\text{b}_{w_{i}}*\mathbf{1}_{{l_{t_{j}}={l_{w_{i}}}}})$ (5)
Scoring method for datasets meeting principle 1 Since the logit as a feature for selection isn’t representative (with negative values), we score the sample $t_{j}$ based on the top-$N_{l}$ linear weighted ranking position score of the words in $S_{r}$ instead of the logit value as Eq 6. We set its ranking position score to zero if $w_{i}$ does not belong to class $l$.
$score_{t_{j}}=\sum_{i=0}^{N_{l}}\left(\frac{2(N_{l}-i)}{(N_{l}+1)N_{l}}*\mathbf{1}_{{l_{t_{j}}={l_{w_{i}}}}}\right)$ (6)
In the k-shot ICL, for each class $l$, the samples with the top-k scores in $D_{l}$ are selected as the demonstration sample for class $l$. The order of the maximum scores in each label determines the demonstration k-shot order.
Class name evaluation To access the semantic information represented by labels in the demonstrations, we analyze logit values of class names in selected samples after selection. We discard the class name $l$ if its logit value, obtained from samples labeled $l$, ranks in the bottom half of $N_{l}$. We then replace it with the label word with top-1 logit to form the initial sample-label pair.
3.4 Multiple-Label Words Insertion in Sample-Label Pairs
@@
We employ multiple label words in sample-label to provide diverse label semantics prompting. Our method sequentially inserts multiple label words into a single sample-label pair, which is more efficient in computation and memory than using multiple augmented sample-label pairs Milios et al. (2023).
The validation set $D_{dev}$ is the train set $D$ with the demonstration samples removed. $D_{dev}^{l}$ is the subset of $D_{dev}$ containing all samples of class $l$. $D_{dev}^{l}$ is evaluated under the initial sample-label pairs obtained from Sec. 3.3 and outputs the logit value of $w$ of class $l$ in $S_{r}$ (excluding words in sample-label pairs). For each label name $l\in L$, we pick the one with the maximum average logit in $D_{dev}^{l}$, and insert it to form sample-multiple-label pairs. This insertion is iterative, with updated demonstrations at every iteration, until no candidates are left in $S_{r}$ or the insertion of the additional label words results in a degraded performance on the validation set.

### v4

# Demonstration Organization via Logit Separability: Advancing In-Context Learning through Multiple Class-Related Words

@@ Abstract
Effective organization of in-context learning (ICL) demonstrations is key to improving the quality of large language model (LLM) responses. To create better demonstrations that guide LLM understanding, we introduce logit separability, a criterion to assess the clarity of both samples and class-related words at the logit level. This facilitates the optimization of sample and label selection, enhancing the precision of information provided in ICL demonstrations. Additionally, we find that incorporating multiple class-related words for each sample, rather than relying on a single class name, improves performance by offering a broader range of label information. Building on these insights, we propose LICL, a logit separability-based method that jointly organizes samples and integrates multiple class-related words into each sample-label pair. Evaluations across seven classification datasets show that this approach significantly improves ICL performance by providing clearer instructions and richer label information [^1] Our code is available at https://github.com/MidiyaZhu/MICL.
Demonstration Organization via Logit Separability: Advancing In-Context Learning through Multiple Class-Related Words
@@ 1 Introduction
In-context learning (ICL) enables large language models (LLMs) to perform new tasks using sample-label pairs as demonstrations, without the need for retraining or fine-tuning (Brown et al., 2020). However, the organization of these sample-label pairs is critical, as it can significantly affect ICL performance [^2] In this paper, we follow Wu et al. (2023) to denote the selection and ranking of sample-label pairs as organization. (Liu et al., 2022).
How do we assess the suitability of samples and labels for demonstration? In traditional machine learning, effective features are those that exhibit high discriminative power, enabling clear distinction between their respective classes (Fukunaga, 2013). Analogously, we introduce logit separability, a criterion to assess how well in-demonstration samples and label words differentiate between classes at the logit level. Logit separability captures two aspects: (1) how significantly a sample is predicted to its true label, where the correct label has substantially higher logit than others, and (2) how consistently a class-related word yields high logit values across samples of the same label. This measure provides insight into both the separability of samples through class-related words and the ability of these words to distinguish between samples of different classes.
(a) Samples’ logit separability in ZSL
(b) Class-related words output distribution in ZSL
(c) Multiple class-related words effectiveness in ICL
@@
Figure 1: Exploration of Samples and class-related words on LLaMA2-7b. (a): Logit separability of SST-2 samples across class-related words under zero-shot learning (ZSL), showing varying degrees of separation due to different input samples. The 1-shot accuracy is demonstrated using the good or bad negative/positive samples, with class names as labels. (b): Logit values of various class-related words for 100 negative and 100 positive SST-2 samples under ZSL, showcasing the logit separability of class-related words across samples. (c): Accuracy comparison in 1-shot ICL using class names, single class-related words, and multiple class-related words (combining the two sets with spaces) as labels. Performance with multiple class-related words surpassed the other two sets. More experiments and analyses, including those on GPT2-xl, are in Appx.A.
Building on this concept, we observe that not all samples or class-related words exhibit the same degree of logit separability. As shown in Fig.1(a), some samples produce strong separation in the logits of class-related words, assigning high logits to words matching the true label and low logits to others, while others show negligible differences between them. Additionally, certain class-related words consistently display higher logit in samples of their corresponding label compared to samples of other labels. As shown in Fig.1(b), "bad" shows a clearer differentiation between positive and negative samples than the class name "negative".
Given the importance of providing clear label-instructed information in demonstrations, leveraging logit separability for selecting samples and labels might enhance ICL performance. Samples with strong logit separability provide more precise signals by aligning class-related words with the model’s internal representations (Fig.1(a)-accuracy). Furthermore, class-related words with distinct logit separability can outperform predefined class names in guiding the model toward accurate predictions. Therefore, selecting both samples and class-related words based on logit separability ensures clearer, more effective demonstrations, ultimately improving the performance of ICL.
Besides, the variability observed in logit values of class-related words across different samples shows that relying on a single class-related word may not adequately capture the full semantics associated with a class name. Unlike expanding candidates via a verbalizer in prediction mapping, directly inserting class-related words into demonstrations offers explicit instructions that guide LLM responses to potential candidates. Consequently, transitioning from single to multiple class-related words could enhance linguistic precision and deepen the semantic representation of class names, potentially leading to enhanced model performance. Fig.1(c) supports that using multiple class-related words in sample-label pairs strengthens ICL performance.
Given samples and labels’ critical roles in demonstration effectiveness for ICL, in this work, we propose a logit separability-based demonstration organization method. We organize demonstration samples based on their zero-shot logit separability across a pool of class-related words. In demonstration label selection, we propose a novel method that utilizes multiple class-related words instead of a single one. The selection, ordering, and quantity of class-related words are determined by the semantic correlations observed in the organized samples’ logit distribution and their validation performance. Moreover, while most research investigates either the sample selection (Zhang et al., 2022; Hongjin et al., 2022; Levy et al., 2023; Lu et al., 2022) or label selection (Yoo et al., 2022; Milios et al., 2023; Gao et al., 2023; Peskine et al., 2023) individually, we comprehensively address both aspects to enhance ICL quality.
@@
1. We introduce Logit Separability, a criterion that evaluates how clearly a sample or label distinguishes between classes at the logit level, helping assess the suitability of samples and labels for demonstrations in ICL.
2. We propose LICL, a Logit separability-based demonstration organization method for ICL. We strategically select and order both samples and labels via the class-related words’ logit separability in LLM’s output space, thereby improving the demonstration’s instruction efficacy.
3. We present a novel label instruction method for ICL that inserts multiple class-related words in sample-label pairs. This approach provides richer label information, improving breadth, reducing ambiguity, and enhancing ICL performance.
@@ 2 Related Work
Label matters in ICL The significance of demonstration labels in ICL has been debated. Yoo et al. (2022) demonstrates a positive relationship between ICL performance and accurate sample-label mapping. Li and Qiu (2023) reveals that using the same sample but different class-related words as labels in the demonstrations can result in very different ICL performance. Wang et al. (2023) suggests that labels derive semantic representations from demonstrations for use in deep layers to make final predictions. Yu and Ananiadou (2024) further shows that these demonstration features are integrated into corresponding labels, with each in-context head extracting features specific to these labels. Milios et al. (2023) uses different class-related words in augmented samples to enhance ICL performance, indicating the potential for enhanced demonstration effectiveness when multiple class-related words are used. However, this approach significantly increases the required token length and running time in augmentation settings.
@@ 3 Method
Figure 2: Overall architecture of LICL: The top part shows the pool refinement, with sample organization based on logit separability across the refined pool. The bottom part presents multiple class-related word insertion via sequential forward search, starting from an initial sample-label pair to form a sample-multiple-label pair.
In this section, we introduce our proposed method LICL, which comprises logit-separability-based demonstrated sample selection and ordering, and corresponding multiple class-related words insertion, as shown in Fig.2.
@@ 3.1 Problem Statement
Give a large language model $M$, label space (class names) $L$, a class-related word pool $P$ with class-related words (includes class names), test sample $x_{test}$ and demonstrations $\left\{x_{i},y_{i}\right\}_{i=1}^{L}$ [^3] Our work focus on 1-shot ICL with label-balanced demonstrations, so labels in demonstration are from the $1$-th to the $L$-th label.. The zero-shot classification of $x_{test}$ can be based on the logit of class names as $\arg\max_{y\in L}p_{M}(y\mid x_{\text{test}})$ or based on the logit of all class-related words in the pool as $\arg\max_{y\in P}p_{M}(y\mid x_{\text{test}})$. Similarly, the 1-shot prediction of $x_{test}$ is $\arg\max_{y\in L}p_{M}(y\mid x_{1}\oplus y_{1},\cdots,x_{L}\oplus y_{L}\oplus x_{\text{test}})$ or $\arg\max_{y\in P}p_{M}(y\mid x_{1}\oplus y_{1},\cdots,x_{L}\oplus y_{L}\oplus x_{\text{test}})$. $\oplus$ is the concatenation operation, and function $p_{M}(\cdot)$ returns the logit of words in $M$’s vocabulary.
3.2 Logit Separability-Based Sample Selection and Ordering
A well-chosen sample should exhibit strong logit separability, producing higher logits for class-related words that correspond to the sample’s label, with these words ranked as highly as possible among all words in the pool $P$ based on logit values. This ensures that the selected samples are aligned with multiple correlated class-related words while minimizing alignment with uncorrelated words, enhancing the sample’s class-related representativeness and improving the effectiveness of ICL demonstrations.
Pool Refinement The initial pool $P$ is constructed from existing works Hu et al. (2022); Zhu et al. (2024) with details in Appx.D. However, not all words in $P$ exhibit clear separable features within both the dataset and the LLM, risking the introduction of incorrect class-related features. To refine $P$ and maintain the quality of the logit separability measurement, we propose a filtering method to generate a refined pool, $P_{r}$, by selecting class-related words based on their logit separability and correlation using point-biserial testing Kornbrot (2014) across dataset samples. Details of this refinement are in Appx.B, with analyses in Appx.H.1.
With the refined pool $P_{r}$ in place, for each zero-shot learning sample, the LLM generates logit values for all words in $P_{r}$. We first remove samples where the class-related word with the highest logit corresponds to an incorrect label. Next, we score samples for selection. Since predictions rely on the word with the highest logit, scoring based on all words introduces noise—particularly in multi-class tasks—because the pool contains far more words linked to incorrect labels than to the correct label. To address this, we focus exclusively on the true label pool, using two methods—Top-Logit Summation and Rank-Weighted Counting. These methods ensure that selected samples align with logit separability across relevant labels, yielding higher logits for class-related words matching the sample’s label and ranking them among the top in $P_{r}$. Let $N_{l}$ be the number of words in $P_{r}$ for class $l$, and $P_{r}^{l}$ be the pool of class-related words for class $l$.
Top-Logit Summation For each training sample $t_{j}$ with label $l_{t_{j}}$, we compute its score by summing the logits of the top $N_{l}$ words from $P_{r}$, sorted in descending order by their logits $b_{w_{i}}$ (Eq.1). Ideally, these top $N_{l}$ words are all class-related words corresponding to $l_{t_{j}}$, indicating strong logit separability. Therefore, only the top $N_{l}$ words $w_{i}$, which belong to the sample’s label $l_{t_{j}}$ (i.e., in $P_{r}^{l_{t_{j}}}$), are included in the sum, ensuring that the score reflects only the contribution from the correct class-related words.
$\text{score}_{t_{j}}=\sum_{i=0}^{N_{l}-1}b_{w_{i}},\text{ if }w_{i}\in P_{r}^{l_{t_{j}}}$ (1)
Rank-Weighted Counting When the logit values of class-related words are negative, summing logit becomes less effective. To address this, we focus on the ranks of the class-related words rather than their logit values to capture logit separability when logit is negative. Each training sample $t_{j}$ with label $l_{t_{j}}$, is evaluated by assigning linear weights to its corresponding class-related words based on their ranks within the top $N_{l}$ words from $P_{r}$, sorted by their logits $b_{w_{i}}$ (Eq.2). Higher-ranked words receive higher weights, emphasizing their prominence.
$\text{score}_{t_{j}}=\sum_{i=0}^{N_{l}-1}\frac{2(N_{l}-i)}{(N_{l}+1)N_{l}}\text{, if }w_{i}\in P_{r}^{l_{t_{j}}}$ (2)
In k-shot ICL, for each class $l$, we select the top-k highest-scoring training samples labeled as $l$. We then construct the demonstration sequence by interleaving these samples across classes: first including the highest-scoring sample from each class, followed by the second-highest, and so on. This ordering ensures that the most logit-separable and representative samples appear early while maintaining balanced class representation, thereby reducing potential bias and enhancing the effectiveness of the demonstrations.
3.3 Multiple Class-Related Words Insertion
As analyzed in Sec.1, relying solely on class names in the sample-label pair may not provide sufficient semantic information for the labels. To address this, we aim to enrich the semantic diversity of label prompts by employing multiple class-related words, proposing the use of a sequence of class-related words to create a sample-multiple-label pair, thereby enhancing the effectiveness of demonstration labels.
@@ 3.3 Multiple Class-Related Words Insertion
After confirming that the selected samples respond well to these multiple class-related words in $P_{r}$, our method sequentially inserts the class-related words into a single sample-label pair to form a sample-multiple-label pair, which is more computationally and memory efficient than using multiple augmented sample-label pairs in ICL instruction.
The word insertion process is based on sequential forward search, where the sample-multiple-label pairs are evaluated as 1-shot demonstrations on the validation set. New class-related words are inserted based on the output logit values. Intuitively, since no demonstration is in place, the labels in the initial sample-label pairs typically use predefined class names. However, we find that predefined class names are not always optimal, as they might exhibit lower logit in zero-shot learning compared to other class-related words. Consequently, we construct the initial sample-label pairs by replacing weak logit response class names with class-related words that yield higher logit values within the selected sample set.
Initial Class-Related Word Updates If a class name $l$ ranks below the top $1$ based on its logit score in zero-shot learning across the selected samples from $P_{r}$, we replace it with the word that has the highest logit value within the selected sample set. This joint process of assigning initial class-related words from the selected samples ensures that the resulting sample-label pairs are better tailored for the LLM’s logit feedback.
Sequential Forward Insertion: After obtaining the initial sample-label pairs, we select and order class-related words to form sample-multiple-label pairs. For each label $l\in L$, we insert the word $w_{i}^{l}$ with the highest average logit value across all samples in $D_{dev}^{l}$ under ICL as Eq.3:
$w_{i}^{l}=\arg\max_{w\in P_{r}^{l}}\frac{1}{|D_{dev}^{l}|}\sum_{x\in D_{dev}^{l}}b_{w}^{x}$ (3)
where $D_{dev}$ is the validation set derived from the training set with the demonstration samples removed, $D_{dev}^{l}$ is the subset of $D_{dev}$ containing all samples of class $l$, and $b_{w}^{x}$ is the logit value of word $w$ for validation sample $x$. After each insertion, the selected word is removed from $P_{r}^{l}$ to prevent repeated insertions.
This insertion is iterative, with the selected word updating sample-multiple-label pairs in each round. The iteration continues until no more candidates remain in $P_{r}^{l}$, or until further insertions reduce validation set performance. Examples of prompts for zero-shot, 1-shot, and 1-shot multiple class-related word learning are provided in Appx.C.